What is the real number property of the following questions?
9(1/9)=1
Commutitative, associative, identity, zero property of multiplication, or multiplication property of -1?
I know it’s not either of the last two.
Thanks!

Found 2 solutions by jim_thompson5910, ikleyn:
Answer by jim_thompson5910(35256)

The commutative property of multiplication is the idea of multiplying two numbers in any order you want. Eg: 2*3 = 3*2 = 6. So we can rule out this property

The associative property of multiplication is a*(b*c) = (a*b)*c which is similar to the commutative property above; but we're using parenthesis to group different pairs of terms. We can rule this out as well.

The identity property of multiplication is the idea of 1 multiplied by any number results in that same number. Example: 7 times 1 = 7. In general, the formula is 1*x = x*1 = x where x is any number. This is close to what we want, but not entirely there. So let's rule this out.

The zero property of multiplication states that multiplying any number by 0 leads to 0. So 0*x = x*0 = 0 regardless of what x is. We can rule this out.

Not sure what you mean by "multiplication property of -1". I think you meant to say "multiplication property of 1" where the 1 has been changed to positive rather than negative. This is restating the "identity property of multiplication". We can rule this out.

Unfortunately there are no other options left. The true answer is "inverse property of multiplication" or "Multiplicative inverse property"

That rule says x*(1/x) = (1/x)*x = 1 where x is any number but zero. Put another way, (a/b)*(b/a) = 1 where a & b are nonzero values. We cannot have zero in the denominator.

Answer by ikleyn(52788)

.

This property is called "multiplicative inverse" or "reciprocal".

It is not listed in your post.


See this Wikipedia article

https://en.wikipedia.org/wiki/Multiplicative_inverse